# Reference Grounded Skill Discovery

### v1

@@ 5.2 Evaluation of discovered skills
Figure 4: Comparison of the learned trajectories of the robot’s base on top-view.
@@
To compare against other methods, we plot the robot’s base trajectories from a top view for the generated skills (Fig. 4), alongside those from the reference motions. The generated trajectories remain clustered around the references, confirming that the discovered skills are consistent, structured variations rather than arbitrary behaviors. Furthermore, note that FID scores for CALM is decrease a lot compare to pure imitation skills. FID for walk becomes 1.4 to 15.5, and becomes 13.9 to 26.7 for run. This means that the policy struggles to learn diverse but at the sametime high fidelity motions. On the other hand, motions produced by RGSD keeps the style since FID score is not much decreased and even for run, FID score increased.
Furthermore, FID scores highlight a key difference. For CALM, FID scores degrade significantly compared to pure imitation: from $1.4\to 15.5$ for walking and from $13.9\to 26.7$ for running, indicating that the policy struggles to generate diverse yet high-fidelity motions. In contrast, motions produced by RGSD preserve the original style, with FID scores remaining stable demonstrating both diversity and fidelity. We attribute this difference to the training setup: in CALM, the policy is always conditioned directly on the motion embedding and therefore never encounters nearby latent vectors during training. By contrast, RGSD encompasses both imitation and discovery phases, where the imitation phase exposes the policy to motion-embedding latents while the discovery phase conditions it on diverse latent vectors. As a result, the policy naturally learns to handle both exact motion embeddings and their local variations.
@@ 5.3 Downstream task evaluation
Figure 5: Training curve
We consider a GoalReaching-Sidestepping task, where the agent must reach a randomly spawned goal within a $20\times 20$ m square arena by sidestepping. We train a high-level policy $\pi_{\text{high}}(z\mid s,g)$ that takes the goal $g$ as input and outputs a high-level action $z$. The learned low-level policy $\pi_{\text{low}}(\cdot\mid s,z)$ then executes the action $a$. During training, we freeze $\pi_{\text{low}}$ and only optimize $\pi_{\text{high}}$. To encourage both goal-reaching and style preservation, we adopt a reward similar to Tessler et al. (2023): $r=\exp(-d_{\text{goal}}^{2})+\exp(-\|z-z_{m}\|^{2}),$ where $d_{\text{goal}}$ is the distance to the goal and $z_{m}$ is the embedding of the sidestepping motion $m$. We conduct training with five different random seeds.
Fig. 5 shows the success ratio of ASE, CALM, and RGSD as training progresses. While CALM and RGSD achieve largely similar success ratios, ASE struggles to reach the goal. We attribute this to the modifications we made to the original ASE implementation, as detailed in Appendix E.2.
To further evaluate motion fidelity, we measured the FID scores between the agent-generated trajectories and the reference sidestepping motion. Initially, ASE appeared to achieve the best score ($20.4$). However, ASE had simply learned to remain standing still to exploit the reward design which is always positive, rather than trying to execute the motion. Between CALM and RGSD, the average FID scores were $46.7$ and $34.3$, respectively, indicating that our method more faithfully reproduces the sidestepping behavior.
5.4 Understanding the challenges of expanding RGSD with METRA
@@
One might suspect that this problem could be bypassed by augmenting the state with additional variables, such as a time variable or global coordinates. However, this does not resolve the issue. In practice, it causes the latent space to be dominated by these added dimensions, leading the policy to focus on them rather than discovering meaningful behaviors. For brevity, we defer a more detailed discussion to Appendix F. To properly represent a substantial portion repetitive motions in dataset, we choose DIAYN as our backbone.
@@ 6 Conclusion
Appendix A From vMF encoder to InfoNCE loss
@@
$\mathcal{L}_{\mathrm{InfoNCE}}=-\log\frac{\exp\!\big(\mathrm{sim}(\mu_{\phi}(s^{a}),z^{+})/T\big)}{\exp\!\big(\mathrm{sim}(\mu_{\phi}(s^{a}),z^{+})/T\big)+\sum_{j=1}^{K}\exp\!\big(\mathrm{sim}(\mu_{\phi}(s^{a}),z_{j}^{-})/T\big)}.$
@@
Appendix B Proof of Within–Motion Alignment
@@
Appendix C Proof of guarantee as an imitation reward
@@
C.1 Proof of reward optimality on motion states
@@
C.2 Proof of local quasi-concavity of the reward
@@ Conclusion
Appendix D Full algorithm of RGSD
@@
Appendix E Experiments details
E.1 State and Action space
We use a 3D SMPL humanoid character, commonly adopted in prior work (Luo et al., 2023; Tessler et al., 2024; Tirinzoni et al., 2025). The state representation is defined as follows:
@@ E.1 State and Action space
E.2 Implementation Details of Baseline algorithms
@@
$\mathcal{J}_{\mathrm{div}}=\mathbb{E}_{d^{\pi}(s)}\;\mathbb{E}_{z_{1},z_{2}\sim p(z)}\left[\left(\frac{D_{\mathrm{KL}}\!\left(\pi(\cdot\mid s,z_{1}),\;\pi(\cdot\mid s,z_{2})\right)}{D_{Z}(z_{1},z_{2})}-1\right)^{2}\right]$
@@
Finally, following conventions in skill discovery frameworks (Gregor et al., 2016; Eysenbach et al., 2018; Laskin et al., 2022; Park et al., 2023b; Rho et al., ), we sampled a skill latent at the beginning of each episode and kept it fixed throughout the rollout. While transitioning between skills is orthogonal and can be incorporated into both our method and the baselines to further improve robustness, we excluded it here to focus on a clean understanding of algorithmic contributions.
E.3 List of reference motions
@@ E.4 Hyperparameters
Optimizer Adam(Kingma & Ba, 2014)
@@
Appendix F Supplementary Discussion: Challenges of Expanding RGSD with METRA
In Section 5.4, we have discussed why the METRA objective conflicts with learning repetitive motions when only local observations are available. A potential workaround is to augment the state with additional variables, such as a global coordinate or time variable. These variables could help distinguish between locally identical motions, e.g., before and after taking a step from walking motion. However, each approach ultimately faces a challenge for different reasons.

### v2

@@ 5.2 Evaluation of discovered skills
Figure 4: Top-view trajectories of the robot’s base when conditioned on latent vectors sampled from the neighborhood of each motion embedding. For each method, we visualize 150 trajectories.
@@
To compare against baselines, we visualize the 150 top-view trajectories of the robot’s base for each method in Fig. 4, alongside the trajectories from the reference motions. Trajectories produced by RGSD remain tightly clustered around the references, indicating that the policy captures the intended motion structure. In contrast, the baseline methods often produce degenerate or drifting trajectories, revealing their inability to preserve the characteristics of the original movements.
These observations are further supported by the FID scores. For three out of four motions, RGSD achieves the lowest FID. While Meta-Motivo typically achieves the second-lowest FID, its generated trajectories still fail to remain centered around the references. CALM shows particularly degraded performance: its FID increases from $1.4$ to $15.5$ for walking and from $13.9$ to $26.7$ for running, indicating difficulty in producing diverse yet high-fidelity behaviors. In contrast, the FID scores of RGSD remain stable, demonstrating that the method successfully preserves motion style while enabling variation.
We attribute this difference to the training setup. In both Meta-Motivo and CALM, the policy is always conditioned directly on the motion embedding and therefore never encounters nearby latent vectors during training. By contrast, RGSD explicitly separates imitation and discovery phases: the imitation phase exposes the policy to motion-embedding latents, while the discovery phase conditions it on diverse samples from the latent neighborhood. Consequently, the policy learns to handle both exact motion embeddings and their local variations, enabling consistent and semantically meaningful skill discovery.
RGSD enables test-time control over behavioral diversity.
A key advantage of RGSD is that it allows users to modulate the diversity of generated behaviors at test time by adjusting the sampling distribution of the latent vector $z$. Recall that we sample $z\sim\mathrm{vMF}(z_{m},\kappa)$, a von Mises–Fisher distribution centered at the motion embedding $z_{m}$ with concentration parameter $\kappa$. A larger $\kappa$ produces samples tightly clustered around $z_{m}$, yielding behaviors that closely imitate the reference motion but exhibit lower diversity. Conversely, a smaller $\kappa$ yields a broader distribution, producing more diverse behaviors that may deviate further from the original motion $m$.
We evaluate this effect on three reference motions by sampling with $\kappa\in\{20,100,1000\}$. Results are shown in Fig. 6. As illustrated, increasing $\kappa$ results in trajectories that stay stylistically similar to the reference motion with limited variation, while decreasing $\kappa$ produces increasingly diverse trajectories that still preserve the core semantic style.
@@ 5.3 Downstream task evaluation
Figure 5: Training curves for the downstream task, along with corresponding FID scores.
Setup.
We consider the GoalReaching task with three style conditions: freestyle, sidestepping, and backward. The agent must reach a randomly spawned goal within a $20\times 20\,\text{m}$ arena while adhering to the specified style. Following the setup in the previous sections, we first train a low-level policy $\pi_{\text{low}}(\cdot\mid s,z)$ using the 20 reference motions for each method. We then train a high-level policy $\pi_{\text{high}}(z\mid s,g)$, which takes the goal $g$ as input and outputs a high-level latent action $z$. The corresponding low-level policy $\pi_{\text{low}}(\cdot\mid s,z)$ then produces the joint action $a$. During this phase, $\pi_{\text{low}}$ is frozen, and we optimize only $\pi_{\text{high}}$.
To encourage both goal reaching and adherence to the commanded style, we adopt a reward similar to tessler2023calm:
$r=\exp(-0.25*\|d_{\text{goal}}\|)+\exp(-8*\|z\oplus z_{m}\|),$
where $d_{\text{goal}}$ is the distance to the goal, and $z\oplus z_{m}$ denotes the cosine distance between the high-level action $z$ and the embedding $z_{m}$ of the directed motion. For the freestyle task, only the first reward term is applied. All experiments are conducted with three random seeds.
Only RGSD consistently respects the commanded style while reaching the goal.
Figure 5 reports the success rates and FID scores of RGSD and baseline methods over the course of training. In the freestyle setting, since no specific style is required, we evaluate only the success rate. Although all methods eventually achieve near-perfect success, RGSD converges significantly faster. In contrast, ASE frequently fails to reach the goal; we observed that it often remains stationary, exploiting the fact that its reward remains positive even without producing the desired motion.
For the sidestepping and backward tasks, we evaluate both the success rate and the FID score. Only RGSD reliably follows the commanded style. Although CALM achieves higher success rates early in training, RGSD eventually matches it. Furtermore, CALM ultimately ignores the style command and simply runs forward. On the other hand, Meta-Motivo respects the style only under certain conditions. For example, in the backward task, it walks backward when the goal is spawned behind the agent, but when the goal appears in front, the agent decides to walk forward. In contrast, RGSD consistently adheres to the commanded style: even when the goal is spawned in front during backward task, it takes a large detour to maintain backward motion. This difference is reflected in the FID scores: Meta-Motivo, CALM, and ASE all exhibit higher FID values, indicating substantial deviations from the intended styles. We refer the reader to the accompanying video for qualitative examples.
Note that these tasks are inherently challenging because the reference motions for sidestepping and backward running do not include any turning behaviors. To reach diverse goal locations while maintaining style fidelity, an agent must discover additional skills that are semantically consistent with the style, such as backward turns at varying angles. Imitation-based baselines lack this flexibility, whereas RGSD acquires a rich set of turning behaviors, enabling it to reach goals while preserving the commanded style.
6 Understanding the challenges of expanding RGSD with METRA
@@
One might suspect that this problem could be bypassed by augmenting the state with additional variables, such as a time variable or global coordinates. However, this does not resolve the issue. In practice, it causes the latent space to be dominated by these added dimensions, leading the policy to focus on them rather than discovering meaningful behaviors. For brevity, we defer a more detailed discussion to Appendix H. To properly represent a substantial portion repetitive motions in dataset, we choose DIAYN as our backbone.
@@ 7 Conclusion
Reproducibility Statement
We have made extensive efforts to ensure the reproducibility of our work across multiple dimensions:
  • A complete pseudo-code description of our algorithm is provided in Appendix F.
  • The full list of reference motions used in our experiments is detailed in Appendix G.3.
  • Comprehensive specifications of states, actions, and hyperparameters are given in Appendix G.4.
  • Rigorous derivations and proofs of our theoretical claims are included in Appendices B, C, and D.1.
  • The full codebase will be publicly released upon acceptance.
Appendix A Qualitative results for controlling diversity experiment
Figure 6: Red dots indicate the agent’s trajectory. RGSD enables test-time control of behavioral diversity by varying the sampling distribution of the latent vector $z$.
Appendix B From vMF encoder to InfoNCE loss
@@
$\mathcal{L}_{\mathrm{InfoNCE}}=-\log\frac{\exp\!\big(\mathrm{sim}(\mu_{\phi}(s^{a}),z^{+})/T\big)}{\exp\!\big(\mathrm{sim}(\mu_{\phi}(s^{a}),z^{+})/T\big)+\sum_{j=1}^{K}\exp\!\big(\mathrm{sim}(\mu_{\phi}(s^{a}),z_{j}^{-})/T\big)}.$ (8)
@@
Appendix C Proof of Within–Motion Alignment
@@
Appendix D Proof of guarantee as an imitation reward
@@
D.1 Proof of reward optimality on motion states
@@
D.2 Proof of local quasi-concavity of the reward
@@ Conclusion
Appendix E Within-Motion Alignment: Latent Space Visualization After Pretraining
Figure 7: Latent embeddings of all 20 reference motions using the pretrained encoder. Each point corresponds to a single frame, and different colors represent different motions.
To evaluate the degree of within-motion alignment, we visualize the latent embeddings of each motion in a 3D space using the pretrained encoder $\mu_{\phi}$. Because the latent vectors are 16-dimensional, we display only their first three dimensions. As shown in Fig. 7, after 1000 epochs of training with the loss in Eq. 8, the state-wise embeddings within each motion form tight clusters and are nearly perfectly aligned, which is consistent with our theoretical results in Section C.
We also quantify this alignment using cosine similarity. For each motion $m_{i}\in\mathcal{M}$, we first compute its motion-level embedding $z_{m_{i}}$ using Eq. 5. Then, for every state $s\in m_{i}$, we measure the cosine similarity between its latent embedding and $z_{m_{i}}$, and average these values to obtain the within-motion alignment score for that motion. Across all 20 motions in the dataset, the average within-motion alignment is $0.99998\pm 8.91\times 10^{-6}$, empirically confirming the effectiveness of the alignment objective.
In practice, overlapping states across different motions can make the motion embeddings harder to distinguish. We mitigate this issue by introducing a phase variable $t$ and concatenating it to each frame $o_{t}$, as well as defining each state as a stack of the most recent $n$ frames, $s_{t}=[o_{t-n+1},o_{t-n+2},\ldots,o_{t}]$. This reduces the likelihood that two motions share identical frame sequences at the same phase $t$. In our dataset, $n=5$ works well, though the optimal value may vary depending on the data. Additionally, our method is robust to a small number of overlapping states, as long as they do not significantly pull the motion-level embeddings of different motions closer together.
Appendix F Full algorithm of RGSD
@@
Appendix G Experiments details
G.1 State and Action space
We use a 3D SMPL humanoid character, commonly adopted in prior work (Luo2023PerpetualHC; tessler2024maskedmimic; tirinzoni2025zero). The state representation is defined as follows:
@@ G.1 State and Action space
G.2 Implementation Details of Baseline algorithms
@@
$\mathcal{J}_{\mathrm{div}}=\mathbb{E}_{d^{\pi}(s)}\;\mathbb{E}_{z_{1},z_{2}\sim p(z)}\left[\left(\frac{D_{\mathrm{KL}}\!\left(\pi(\cdot\mid s,z_{1}),\;\pi(\cdot\mid s,z_{2})\right)}{D_{Z}(z_{1},z_{2})}-1\right)^{2}\right]$ (13)
@@
Finally, following conventions in skill discovery frameworks (vic; diayn; cic; metra; rholanguage), we sampled a skill latent at the beginning of each episode and kept it fixed throughout the rollout. While transitioning between skills is orthogonal and can be incorporated into both our method and the baselines to further improve robustness, we excluded it here to focus on a clean understanding of algorithmic contributions.
G.3 List of reference motions
@@ G.4 Hyperparameters
Optimizer Adam(kingma2014adam)
@@
Appendix H Supplementary Discussion: Challenges of Expanding RGSD with METRA
In Section 6, we have discussed why the METRA objective conflicts with learning repetitive motions when only local observations are available. A potential workaround is to augment the state with additional variables, such as a global coordinate or time variable. These variables could help distinguish between locally identical motions, e.g., before and after taking a step from walking motion. However, each approach ultimately faces a challenge for different reasons.